OpenStudy (anonymous):

the 3 should be on top of the sqrt

OpenStudy (anonymous):

can you draw it out

OpenStudy (anonymous):



OpenStudy (anonymous):

wouldnt the answer be 3-x?

OpenStudy (phi):

It's not much simpler, but you could rewrite this as \[\sqrt[3]{81x}-\sqrt[3]{3x^{4}}= 3\cdot \sqrt[3]{3x}-x\cdot \sqrt[3]{3x}\] \[= (3-x) \cdot \sqrt[3]{3x}\]

OpenStudy (anonymous):

ohh theres more to the problem? well the begging is 3-x i agree.

OpenStudy (anonymous):

beggining**

OpenStudy (anonymous):

it would be (3-x)(3-x), which is x^2-6x+9.....